What is the value today of a money machine that will pay $2,756.00 per year for 30.00 years? Assume the first payment is made 7.00 years from today and the interest rate is 6.00%.

What is the value today of a money machine that will pay $1,720.00 every six months for 29.00 years? Assume the first payment is made six months from today and the interest rate is 6.00%.

What is the value today of a money machine that will pay $1,060.00 per year for 10.00 years? Assume the first payment is made today and that there are 10.0 total payments. The interest rate is 7.00%.

Derek will deposit $3,988.00 per year for 19.00 years into an account that earns 6.00%. The first deposit is made next year. How much will be in the account 19.0 years from today?
Answer #1

1. Value of annuity 7 years from now using annuity formula = PMT*(1-(1+r)-n)/r = 2756*(1-(1+6%)-30)/6% = 37,935.87
PV of annuity today = 37,935.87/(1+6%)7 = 25,229.52

2. Rate half yearly = 6%/2 = 3%
Number of Periods = 29*2 = 58
PV of annuity today = PMT*(1-(1+r)-n)/r = 1720*(1-(1+6%)-58)/6% =27,690.25

3. PV of annuity due = (1+r)*PMT*(1-(1+r)-n)/r = (1+6%)*1060*(1-(1+6%)-58)/6% = 7801.13

4. FV of annuity = PMT*((1+r)n-1)/r = 3988*((1+6%)19-1)/6% = 134,634.85

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well